OpenStudy (anonymous):

ok i have f(x)=x-2, calculate f(-3). no idea what im doing S.O.S.

OpenStudy (anonymous):

f(-3) = -3 -2 = -5

OpenStudy (anonymous):

plug -3 in place of x

OpenStudy (anonymous):

Well, calulating f(-3) means insert -3 in place of f(x).

OpenStudy (anonymous):

oh ok, that helps!! thanks

OpenStudy (anonymous):

so what do i do with the f?

OpenStudy (anonymous):

The f is simply a label, it's not actually part of the equation. If you are told 'Mix color red with color blue to make purple.', mathematically, it would be 'Purple=color(red)+color(blue)'. Does that help you understand?

OpenStudy (anonymous):

ya thanks!

OpenStudy (anonymous):

Couldn't have explained it better than sharktopi....

OpenStudy (anonymous):

great explanation

OpenStudy (anonymous):

so then getting replacing the x's and f's the problem would look like this -3=-3-2?

OpenStudy (anonymous):

No, it would be f(x)=-3-2. f(x) is just code for 'y' :)

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

then what? sorry new at pre alg.

OpenStudy (anonymous):

It's cool, I'm barely doing Algebra 2 :) f(x)=-3-2 Remember, f(x) is code for 'y', therefore y=-3-2 Sicne -3-2 is -5 y=-5

OpenStudy (anonymous):

oh ok know i get it! thanks a lot!